OpenStudy (anonymous):

integrate from 1 to 0 for 2/ (1+x^2) dx

OpenStudy (unklerhaukus):

\[\int\limits_1^0\frac{2}{1+x^2}\text dx\] did you say the limits in the right order

OpenStudy (anonymous):

yeah i did its from 0 to 1

OpenStudy (unklerhaukus):

\[\int\limits_0^1\frac{2}{1+x^2}\text dx\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

\[=2\int\limits_{0}^{1} 1/(1+x^2) dx\]\[=[2\tan^{-1} x]_{0}^{1}\]\[=2\tan^{-1} 1 - 2\tan^{-1} 0\]

OpenStudy (anonymous):

\[=2(\pi/4) -0\]\[=\pi/2\]

OpenStudy (anonymous):

why is it tan?

OpenStudy (lalaly):

http://www.maa.org/pubs/Calc_articles/ma024.pdf

OpenStudy (anonymous):

the integration of 1/(1+x^2) with respect to x is tan^-1 x not tan x.